# Rethinking Oversmoothing in Graph Neural Networks: A Rank-Based Perspective

### v3

@@ A.3 An illustrative example
A.6 proof of Proposition 5.5
@@ Appendix B Limitations
Although a small effective rank or numerical rank indicates oversmoothing and can be subsequently linked to the underperformance of GNNs, a large effective rank or numerical rank does not necessarily correspond to a good network performance. Prior study has suggested some degree of smoothing can be beneficial [16], and as an extreme example, features sampled from a uniform distribution and with randomly assigned labels would almost surely have a large effective rank, but they cannot be classified accurately due to the lack of any underlying pattern. We note that this limitation is not specific to the two relaxed rank measures, as the same argument is directly applicable to all other oversmoothing metrics.
Consequently, as shown in Table 4, when additional components are used to (partially) alleviate oversmoothing, particularly when residual connections are used, the accuracy ratio may remain large over the layers, and all oversmoothing metrics correlate poorly with the accuracy of GNNs. This in-turn suggests these oversmoothing metrics become less informative as the oversmoothing problem is mitigated or alleviated.
@@ Appendix C Computational Complexity Analysis
In general, the computational cost of metrics to quantify oversmoothing is marginal compared to the cost of training. In production, measuring the emergence of oversmoothing is done by training the model and checking the performance on the validation and test sets as the number of layers changes. The cost of additionally computing effective or numerical ranks is marginal. Moreover, we note that computing the metrics on a subsection of a graph can be sufficiently informative, and as a consequence, most oversmoothing metrics studied in this paper can be computed in less than $10~{}ms$ for a graph (or a subsection of it) with less than 2000 nodes.
@@ Appendix D Additional Experimental Results with Synthetic Weights
  • to demonstrate that similar untrained asymptotic experiments are inherently unrealistic, despite being extensively used in the literature [44, 35, 34, 45, 29, 42], as they fail to reliably capture oversmoothing in shallower GNNs where the performance degradation occurs.
  • to examine the convergence properties of different oversmoothing metrics with weight size control and to empirically validate Theorems 5.1 and 5.4.
@@
In this asymptotic synthetic setting, as presented in Table 2, the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ exhibit decay patterns similar to those of the effective rank and numerical rank, suggesting these metrics are equally sensitive to asymptotic rank collapse. However, this behaviour stands in stark contrast to results on trained networks presented in Tables 1 and 3 and Appendix G, where the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ often fail to detect oversmoothing.
Moreover, these asymptotic results validate Theorems 5.1 and 5.4, showing that the numerical rank converges to one for GCN + LReLU and GAT + any subhomogeneous activation functions. Without making any additional assumption on the normalization of the adjacency matrix, the effective rank and numerical rank do not generally decay to one when subhomogeneous activation functions, e.g. Tanh, are used in GCNs.
@@ Appendix E Additional Experimental Results on Activation Functions and Different Datasets
We extend Table 1 to subhomogenous Tanh activation function and a few additional datasets. The experimental setting is consistent with that of Section 6.
@@ Appendix F Additional Experimental Results with Different Network Components
Prior literature indicates that when adding additional components, such as bias [33] or residual terms [36], the Dirichlet energy does not decay. Therefore, we compare the correlation coefficients in Table 4 between existing oversmoothing metrics when additional components are added, such as bias, LayerNorm, BatchNorm, PairNorm [46], DropEdge [28, 15] and residual connections [36].
All experiments follow the setup described in Section 6. In addition, DropEdge has a probability of 0.5 in removing each edge at each layer. The residual connection is implemented as follows
$X^{(l+1)}=\sigma(AX^{(l)}W_{1}^{(l)})+X^{(0)}W_{2}^{(l)}$
Table 4 demonstrates that both the effective rank and numerical rank achieve a higher average correlation with the classification accuracy than alternative metrics. This finding confirms their superior consistency in detecting oversmoothing across a variety of architectural variants.
Furthermore, we note that when oversmoothing is effectively alleviated, e.g. when residual connection is used, all metrics have a poor correlation with the classification accuracy. This generic limitation is discussed in Appendix B.

### v4

@@ A.1 Equivalence of $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$
A.6 Some situations where $\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0$
In this section we explore three different situations where the Hilbert distance of the features from the dominant eigenvector $u$ of the thea aggregation matrices is guaranteed to converge to $1$.
1st situation
A first situation where the Hilbert distance of the features from the dominant eigenvector $u$ converges to zero is the case of all matrices $A^{(l)}$ are contractive.
Lemma A.2.
Let $A^{(l)}$ be nonnegative and irreducible matrices with dominant eigenvector $u\in\mathcal{K}$. Assume also $X^{(0)}$ to be strictly positive, $W^{(l)}$ nonnegative with $\min_{j}\max_{i}W_{ij}^{(l)}>0$ and $\sigma\in C(\mathbb{R}^{N}_{+},\mathbb{R}^{N}_{+})$ a nonlinear function that is order preserving, subhomogeneous and such that $u$ is also an eigenvector of $\sigma$. Any matrix $A^{(l)}$ is known to satisfy $d_{H}(A^{(l)}x,A^{(l)}y)\leq d_{H}(x,y)$ for all $x,y\in\mathcal{K}$, then if $\lim_{k\rightarrow\infty}\prod_{l=1}^{k}\beta_{l}=0$
$\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0.$ (41)
Proof.
The proof is a trivial consequence of Lemma˜5.2. Indeed iterating the result in the thesis we know that
$\max_{i}d_{H}\left(\sigma\big(X^{(l)}_{:,i}\big),u\right)\leq\prod_{i=1}^{l-1}\beta_{i}\max_{i}d_{H}(X_{:,i}^{(0)},u),$ (42)
concluding the proof ∎
In particular we remind that from the Perron-Frobenius theory any strictly positive matrix $A$ is known to be contractive of a parameter $\beta<1$, see lemmensNonlinearPerronFrobeniusTheory2012.
2nd situation
A second situation where the Hilbert distance of the features from the dominant eigenvector $u$ converges to zero is the case of a strictly subhomogeneous activation function.
Lemma A.3.
Let $A^{(l)}$ be nonnegative and irreducible matrices with dominant eigenvector $u\in\mathcal{K}$. Assume also $X^{(0)}$ to be strictly positive, $W^{(l)}$ nonnegative with $\min_{j}\max_{i}W_{ij}^{(l)}>0$ and $\sigma\in C(\mathbb{R}^{N}_{+},\mathbb{R}^{N}_{+})$ a nonlinear function that is order preserving, strongly subhomogeneous and such that $u$ is also an eigenvector of $\sigma$. Since $\sigma$ is strongly subhomogeneous, for any $l$ there exists $\beta_{l}<1$ such that $\max_{i}d_{H}(\sigma(A^{(l)}X^{(l)}W^{(l)})_{:,i},u)\leq\beta_{l}d_{H}((A^{(l-1)}X^{(l-1)}W^{(l-1)})_{:,i},u)$, then if $\lim_{k\rightarrow\infty}\prod_{l=1}^{k}\beta_{l}=0$
$\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0.$ (43)
Proof.
The proof is again a trivial consequence of Lemma˜5.2. Indeed iterating the result in the thesis we know that
$\max_{i}d_{H}\left(\sigma\big(X^{(l)}_{:,i}\big),u\right)\leq\prod_{i=1}^{l-1}\beta_{i}\max_{i}d_{H}(X_{:,i}^{(0)},u),$ (44)
concluding the proof. ∎
Note that whenever we have a strongly concave activation function $\sigma$ on $\mathbb{R}_{+}$, e.g. tanh, then it is strongly subhomogeneous and so the result above applies.
3rd situation
Here we discuss a third situation, possibly weaker than the previous ones. This is essentially an adaptation of the ergodic theorem proved in Nussbaum (1990). Consider the cone $\mathcal{K}^{\prime}=\mathbb{R}_{+}^{N\times d}$ and introduce the set $\Pi=\{Y\in\mathcal{K}^{\prime}\text{ s.t. }Y_{:,i}=\alpha_{i}u\;\forall i=1,\dots,d\}$. Moreover let $\psi\in\mathrm{Int}(\mathcal{K}^{\prime})$ and define $\Sigma_{\psi}=\{Y\in\mathcal{K}^{\prime}\text{ s.t. }\psi(Y)=1\}$. Finally given $X\in\mathcal{K}^{\prime}$ define
$\pi(X):=\operatorname*{arg\,min}_{Y\in\Sigma_{\psi}}d_{H}(X,Y).$ (45)
Definition A.4 (Hypotheses H).
We say that a GNN $X^{(l+1)}=f^{(l)}(X^{(l)})=\sigma(A^{(l)}X^{(l)}W^{(l)})$ with $X^{(0)}\in\mathcal{K}^{\prime}$ satisfies hypotheses $[H]$ if the following conditions are verified:
  1. H1
    $A^{(l)}$ is nonnegative and $\min_{j}\max_{i}W^{(l)}_{ij}>0$ for any $l$.
  2. H2
    The function $\sigma$ is subhomogeneous and differentiable in $\mathbb{R}_{+}$.
  3. H3
    $u\in\mathcal{K}$ is the dominant eigenvector of all of the matrices $A^{(l)}$ and it is also an eigenvector of $\sigma$.
  4. H4
    There exists an integer $p>0$ and a sequence of $dN\times dN$ strictly positive matrices $\{B^{(l)}\}$ such that $\forall X$ with $m\big(X^{(lp)},\pi(X^{(lp)})\big)\pi(X^{(lp)})\leq X\leq M\big(X^{(lp)},\pi(X^{(lp)})\big)\pi(X^{(lp)})$
    $\partial_{X}g^{(l)}(X)\geq B^{(l)}\quad\forall l\geq 1$
    where $g^{(l)}(X):=f^{((l+1)p-1)}\circ\dots\circ f^{(lp)}(X)$.
  5. H5
    $\forall l\geq 0$ there exists $\eta^{(l)}>0$ s.t. $B^{(l)}\pi(X^{(lp)})\geq\eta^{(l)}g^{(l)}\big(\pi(X^{(lp)})\big)$.
  6. H6
    $\lim_{M\rightarrow\infty}\sum_{l=0}^{M}\eta^{(l)}exp\big(-\Delta(B^{(l)})\big)=\infty$ where $\Delta(B)=\sup_{x,y\in\mathcal{K}^{\prime}}d_{H}(Bx,By)<\infty$ is the projective diameter of the matrix $B$.
Theorem A.5.
Let $X^{(l+1)}=f^{(l)}(X^{(l)})=\sigma(A^{(l)}X^{(l)}W^{(l)})$, with $X^{(0)}\in\mathcal{K}^{\prime}$, be a GNN satisfying hypotheses $[H]$. Then
$\lim_{l\rightarrow\infty}\max_{i}\big(d_{H}(X^{(l)}_{:,i},u)\big)=0.$
Proof.
We claim that under hypotheses [H]
$\lim_{l\rightarrow\infty}d_{H}(X^{(l)},\Pi)=0.$ (46)
As a consequence of this it is easy to note that $\lim_{l\rightarrow\infty}\max_{i}\big(d_{H}(X^{(l)}_{:,i},u)\big)=0$. Indeed it is not difficult to check that
$exp\Big(d_{H}\big(X^{(l)},\pi\big(X^{(l)}\big)\Big)$ $=\frac{M(X^{(l)},\pi\big(X^{(l)}\big)}{m(X^{(l)},\pi\big(X^{(l)}\big)}=\max_{ji}\max_{hk}\frac{X^{(l)}_{ji}\pi\big(X^{(l)}\big)_{hk}}{X^{(l)}_{hk}\pi\big(X^{(l)}\big)_{ij}}\geq$ (47)
$\geq\max_{i}\max_{j}\max_{h}\frac{X^{(l)}_{ji}\pi\big(X^{(l)}\big)_{hi}}{X^{(l)}_{hi}\pi\big(X^{(l)}\big)_{ji}}=\max_{i}exp\Big(d_{H}(X^{(l)}_{:,i},u)\Big).$
where we have used that by definition $\pi\big(X^{(l)}\big)_{:,i}=\alpha_{i}u$ for all $i$, where necessarily $\alpha_{i}>0$ for all $i$. Otherwise, since $X^{(l)}\in\mathcal{K}^{\prime}$, we would have $d_{H}\big(X^{(l)},\pi\big(X^{(l)})\big)=\infty$, against the minimality.
Next we prove the claim. The proof is adapted for our scopes from the proof of the weak ergodic theorem 2.1 proved in Nussbaum (1990) for homogeneous mappings. To simplify the notation we denote by $X:=X^{(lp)}$, $\pi:=\pi(X^{(lp)})$, $g:=g^{(l)}$, $m:=m(X^{(lp)},\pi(X^{(lp)}))$, $B:=B^{(l)}$, $\eta:=\eta^{(l)}$ and $M:=M(X^{(lp)},\pi(X^{(lp)}))$. Then consider $z^{1}(t)=(1-t)m\pi+tX$ and $z^{2}(t)=(1-t)X+tM\pi$. Then from hypothesis [H4] we have the following inequalities:
$g(X)-g(m\pi)=\int_{0}^{1}\partial_{X}g(z^{1}(t))\big(X-m\pi\big)\geq B\big(X-m\pi\big)$ (48)
$g(M\pi)-g(x)=\int_{0}^{1}\partial_{X}g(z^{2}(t))\big(M\pi-X\big)\geq B\big(M\pi-X\big).$
In particular, since $g$ is subhomogeneous and by definition of $\pi$ $m\leq 1$ and $M\geq 1$ we have:
$mg(\pi)+B(X-m\pi)$ $\leq g(m\pi)+B(X-m\pi)\leq g(X)\leq$ (49)
$\leq g(M\pi)-B(M\pi-X)\leq Mg(\pi)-B(M\pi-X)$
Since $B(X-m\pi)+B(M\pi-X)=(M-m)B\pi$, we can use Lemma 2.2 in Nussbaum (1990) we know that
$B(X-m\pi)\geq\gamma(M-m)B\pi\qquad\text{or}\qquad B(M\pi-X)\geq\gamma(M-m)B\pi,$ (50)
where $\gamma=(1/2)exp(-\Delta(B))$. Without loss of generality assume that $B(X-m\pi)\geq\gamma(M-m)B\pi$ the second case can be handled analogously. Then, since $B(M\pi-X)$, from (51), we have
$mg(\pi)+\gamma\eta(M-m)g(\pi)$ $\leq mg(\pi)+\gamma(M-m)B\pi\leq g(X)\leq Mg(\pi).$ (51)
In particular
$d_{H}(g(X),g(\pi))\leq\log(\frac{M}{m+\eta\gamma(M-m})$ $=\log\Big(\frac{M}{m+\eta\gamma(M-m)}\Big)=$ (52)
$=\log(M/m)\frac{\log\Big(\frac{M}{m}\frac{1}{1+\eta\gamma(M/m-1)}\Big)}{\log(M/m)}\leq$
$\leq\log(M/m)(1-\eta\gamma)=(1-\eta\gamma)d_{H}(X,\pi).$
In the last we have used the following fact: if $\xi_{1}(s)=\log\big(s(1+\eta\gamma(s-1))^{-1}\big)$ and $\xi_{2}(s)=\log(s)$ with $s>1$ then since $\xi_{1,2}(1)=0$ and $1+\eta\gamma(s^{\prime}-1)>1$ for $s^{\prime}>1$, for any $s>1$ there exists some $s^{\prime}>1$ such that
$\xi_{1}(s)/\xi_{2}(s)=\xi^{\prime}_{1}(s^{\prime})/\xi^{\prime}_{2}(s^{\prime})=(1-\eta\gamma)(1+\eta\gamma(s^{\prime}-1))^{-1}\leq(1-\eta\gamma).$ (53)
Now note that $d_{H}(X,\pi)=d_{H}(X,\Pi)$ by the minimality of $\pi$ and the fact that $d_{H}(X,\alpha Y)=d_{H}(X,Y)$ for any $\alpha>0$. Second, using [H1] and [H3] it is very easy to observe that $g(\pi)\in\Pi$. So $d_{H}(g(X),\Pi)\leq d_{H}(g(X),g(\pi))$. In conclusion we have proved that
$d_{H}\big(X^{(lp)},\Pi\big)=d_{H}\big(g^{(l)}(X^{(l)}\big),\Pi)\leq\left(1-\eta^{(l)}\frac{exp\big(-\Delta(B^{(l)}\big)}{2}\right)d_{H}(X^{(l)},\Pi).$ (54)
In particular iterating, using (47) and recalling that $\max_{i}d_{H}(X^{(l_{1})}_{i},u)\leq\max_{i}d_{H}(X^{(l_{2})}_{i},u)$ if $l_{2}>l_{1}$ (see Lemma˜5.2) we have that for any $L>lp$:
$\max_{i}d_{H}(X^{(L)}_{i},u)$ $\leq\max_{i}d_{H}(X^{(lp)}_{i},u)\leq d_{H}\big(X^{(lp)}_{i},\Pi\big)\leq$ (55)
$\leq\prod_{j=0}^{l}\left(1-\eta^{(j)}\frac{exp\big(-\Delta(B^{(j)}\big)}{2}\right)d_{H}(X^{(0)},\Pi).$
Moreover, we have that
$\lim_{l\rightarrow\infty}\prod_{j=0}^{l}\left(1-\eta^{(j)}\frac{exp\big(-\Delta(B^{(j)}\big)}{2}\right)$ $=0\iff$ (56)
$\lim_{l\rightarrow\infty}\sum_{j=0}^{l}-\log\left(1-\eta^{(j)}\frac{exp\big(-\Delta(B^{(j)}\big)}{2}\right)$ $=\infty\iff$
$\lim_{l\rightarrow\infty}\sum_{j=0}^{l}\left(\eta^{(j)}\frac{exp\big(-\Delta(B^{(j)}\big)}{2}\right)$ $=\infty,$
which concludes the proof.
Next, we discuss briefly the hypothesis H4 which is the most technical one. In particular, such hypothesis recalls the property of a primitive matrix. Actually we note that, if the matrices $A^{(l)}$ are primitive, under few additional mild assumptions, it is possible to prove that the hypothesis H4 is always satisfied. Assume that there exists an index $p$ such that:
  1. 1.
    The matrices $A^{(l)}$ with $l=0,\dots,p-1$ satisfy $A^{(l)}\geq\alpha_{1}A^{*}$ where $A^{*}$ is the unweighted adjacency matrix of the graph; i.e. $A^{*}_{ij}=1$ if the edge $(i,j)$ belongs to the graph, $A^{*}_{ij}=0$ otherwise. And assume that $A^{*}$ is primitive of index smaller then $p$, i.e. $\big((A^{*})^{p}\big)_{ij}>1$ for all $i,j$
  2. 2.
    $\partial_{x}\sigma(X_{ij})\geq\alpha_{2}$ for all $X$ with $X=f^{(l)}\circ\dots\circ f^{(0)}(X^{*})$ varying $l=0,\dots,p-1$ and $X^{*}$ among the matrices that satisfy $m\big(X^{(0)},\pi(X^{(0)}\big)\pi(X^{(0)})\leq X\leq M\big(X^{(0)},\pi(X^{(0)})\big)\pi(X^{(0)})$.
  3. 3.
    $\big(W^{(0)}\cdot\dots\cdot W^{(p-1)}\big)_{ij}>\alpha_{3}>0$ for all $i,j$.
Then for any matrix $X$ satisfying $m\big(X^{(0)},\pi(X^{(0)})\big)\pi(X^{(0)})\leq X\leq M\big(X^{(0)},\pi(X^{(0)})\big)\pi(X^{(0)})$ we have
$\Big(\partial_{X}g(X)\Big)_{ij}\geq\alpha_{1}^{p}\alpha_{2}^{p}\alpha_{3}\quad\forall ij$
where $g(X):=f^{(p-1)}\circ\dots\circ f^{(0)}(X)$.
Indeed, vectorizing the matrix $X$ we can write $\partial_{X}f^{(i)}(X)=\mathrm{diag}\Big(\sigma^{\prime}\big((W^{T}\otimes A)vec(X)\big)\Big)(W^{T}\otimes A)\geq\alpha_{2}(W^{T}\otimes A).$ Then we can apply the chain rule to the function $g$ to get
$\Big(\partial_{X}g(X^{*})\Big)\geq\alpha_{2}^{p}\big({W^{(p-1)}}^{T}\otimes A^{(p-1)}\big)\dots\big({W^{(0)}}^{T}\otimes A^{(0)}\big).$ (57)
Finally we use the properties of the matrices $W^{(i)}$ and $A^{(i)}$ and the properties of the Kronecker product to obtain
$\Big(\partial_{X}g(X^{*})\Big)\geq\alpha_{2}^{p}\big({W^{(p-1)}}^{T}\dots W^{(0)}{}^{T}\big)\otimes\big(A^{(p-1)}\dots A^{(0)}\big)\geq\alpha_{2}^{p}\alpha_{1}^{p}\alpha_{3}\mathds{1}$ (58)
where $\mathds{1}$ is the matrix with every entry equal to $1$.
A.7 Eigenpairs of entry-wise subhomogeneous maps
We conclude with a formal investigation of the eigenpairs of activation functions that are entry-wise subhomogeneous. Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R},\mathbb{R})$ that is subhomogeneous on $\mathbb{R}_{+}$. Then one can easily show that $\sigma$ is itself subhomogeneous on $\mathbb{R}^{N}_{+}$. We have the following result,
Proposition A.6.
Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R}_{+},\mathbb{R}_{+})$ be order preserving. Then: 1) If $\sigma$ is homogeneous, any positive vector is an eigenvector of $\sigma$, 2) If $\sigma$ is strictly subhomogeneous, the only eigenvector of $\sigma$ in $\mathcal{K}$ is the constant vector.
@@ Appendix B Limitations
Although a small effective rank or numerical rank indicates oversmoothing and can be subsequently linked to the underperformance of GNNs, a large effective rank or numerical rank does not necessarily correspond to a good network performance. Prior study has suggested some degree of smoothing can be beneficial Keriven (2022), and as an extreme example, features sampled from a uniform distribution and with randomly assigned labels would almost surely have a large effective rank, but they cannot be classified accurately due to the lack of any underlying pattern. We note that this limitation is not specific to the two relaxed rank measures, as the same argument is directly applicable to all other oversmoothing metrics.
Consequently, as shown in Table˜4, when additional components are used to (partially) alleviate oversmoothing, particularly when residual connections are used, the accuracy ratio may remain large over the layers, and all oversmoothing metrics correlate poorly with the accuracy of GNNs. This in turn suggests these oversmoothing metrics become less informative as the oversmoothing problem is mitigated or alleviated.
@@ Appendix C Computational Complexity Analysis
In general, the computational cost of metrics to quantify oversmoothing is marginal compared to the cost of training. In production, measuring the emergence of oversmoothing is done by training the model and checking the performance on the validation and test sets as the number of layers changes. The cost of additionally computing effective or numerical ranks is marginal. Moreover, we note that computing the metrics on a subsection of a graph can be sufficiently informative, and as a consequence, most oversmoothing metrics studied in this paper can be computed in less than 10ms10~ms for a graph (or a subsection of it) with less than 2000 nodes.
@@ Appendix D Additional Experimental Results with Synthetic Weights
  • to demonstrate that similar untrained asymptotic experiments are inherently unrealistic, despite being extensively used in the literature Wang et al. (2022); Rusch et al. (2022; 2023b); Wu et al. (2023); Roth (2024); Wang et al. (2025), as they fail to reliably capture oversmoothing in shallower GNNs where the performance degradation occurs.
  • to examine the convergence properties of different oversmoothing metrics with weight size control and to empirically validate Theorems˜5.1 and 5.3.
@@
In this asymptotic synthetic setting, as presented in Table˜2, the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ exhibit decay patterns similar to those of the effective rank and numerical rank, suggesting these metrics are equally sensitive to asymptotic rank collapse. However, this behaviour stands in stark contrast to results on trained networks presented in Tables˜1 and 3 and Appendix˜G, where the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ often fail to detect oversmoothing.
Moreover, these asymptotic results validate Theorems˜5.1 and 5.3, showing that the numerical rank converges to one for GCN + LReLU and GAT + any subhomogeneous activation functions. Without making any additional assumption on the normalization of the adjacency matrix, the effective rank and numerical rank do not generally decay to one when subhomogeneous activation functions, e.g. Tanh, are used in GCNs.
@@ Appendix E Additional Experimental Results on Activation Functions and Different Datasets
We extend Table˜1 to subhomogenous Tanh activation function and a few additional datasets. The experimental setting is consistent with that of Section˜6.
@@ Appendix F Additional Experimental Results with Different Network Components
Prior literature indicates that when adding additional components, such as bias Rusch et al. (2023a) or residual terms Scholkemper et al. (2024), the Dirichlet energy does not decay. Therefore, we compare the correlation coefficients in Table˜4 between existing oversmoothing metrics when additional components are added, such as bias, LayerNorm, BatchNorm, PairNorm Zhao and Akoglu (2019), DropEdge Rong et al. (2019); Huang et al. (2020) and residual connections Scholkemper et al. (2024).
All experiments follow the setup described in Section˜6. In addition, DropEdge has a probability of 0.5 in removing each edge at each layer. The residual connection is implemented as follows
$X^{(l+1)}=\sigma(AX^{(l)}W_{1}^{(l)})+X^{(0)}W_{2}^{(l)}.$
Table˜4 demonstrates that both the effective rank and numerical rank achieve a higher average correlation with the classification accuracy than alternative metrics. This finding confirms their superior consistency in detecting oversmoothing across a variety of architectural variants.
Furthermore, we note that when oversmoothing is effectively alleviated, e.g. when residual connection is used, all metrics have a poor correlation with the classification accuracy. This generic limitation is discussed in Appendix˜B.